OpenStudy (anonymous):

The price of a cup of coffee was $2.15 yesterday. Today, the price rose to $2.50 . Find the percentage increase. Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

(2.50-2.15)*100 ---------- 2.15

OpenStudy (anonymous):

2.15 - 2.50 = -0.35 -0.35 ----- <--- division bar 2.15 = -0.1627... (100 x -0.1627...)% = -16.3 change the negative to a positive. 16.3 is your answer